nou. I zanat sane lomo s* 1515906654 l i d , where & = {(1,4,7) 1 Sys 4 y 5:540SS }.
Answer #1

リニ4224=2 ædady = ( 72 72 dxdzdy x²+22 Y = 1 = y n=0 y=42=4 9-12=4 17. y=42=4 (tan) -ton.co) dzdy y = 7 Z-y 4 4 2=4 (-o) dzdy